OpenStudy (anonymous):

f(x)=x^2+3x; g(x)=x+3 Find (f o g)(5)

OpenStudy (anonymous):

EDIT: (f o g) means plug function g into function f, so take the function g and plug that function in wherever you see a 'x' in function f. So f=x^2+3x becomes f=(x+3)^2+3(x+3). Then you just have to plug in 5 wherever you see a x. So, plugging in 5: f= (5+3)^2+3(5+3)= 64+24=88